Calculate the freezing point depression expected for 0.0755 m aqueous solution of Na2SO4. If this solution actually freezes at –0.480°C, what would be the value of van’t Hoff factor? (Kf for water = 1.75 K kg/mol).

Grade:12

1 Answers

Patel priyanka
6 years ago
Dispersion in freezing point, expected∆Tf=T°-Tf=Kfmm=0.0755m, kf=1.75 K kg/mol ,∆Tf=1.75×0.0755=0.132℃If TF=-0.480℃Then ,∆Tf=0.480℃=iKfm 0.480 =i1.75x0.0755 i=1.75×0.0755÷0.480 i=0.00027